(2-48) Promotion of a Rock Concert

BBT Productions, Ltd., is promoting a rock concert in London. The bands will receive a flat

fee of £7 million in cash. The concert will be shown worldwide on closed-circuit television.

BBT will collect 100% of the receipts and will return 30% to the individual local

closed-circuit theater managers. BBT expects to sell 1.1 million seats at a net average

price of £13 each. BBT will also receive £300,000 from the London arena (which has sold

out its 19,500 seats, ranging from £150 for box seats to £20 for general admission, for a

gross revenue of £1.25 million); BBT will not share the £300,000 with the local promoters.

1) The general manager of BBT Productions is trying to decide what amount to spend

for advertising. What is the most BBT could spend and still break even on overall

operations, assuming sales of 1.1 million tickets?

2) If BBT desires an operating income of £500,000, how many seats would it have to sell?

Assume that the average price is £13 and the total fixed costs (including £2,000,000

in advertising) are £9 million.
Answer #1
1) A Number of tickets sold               1,100,000
B Collection per ticket £                   13.00
C=A*B Total collection from ticke sale £        14,300,000
D Receipt from London Arena £              300,000
E=C+D Total Receipts £           14,600,000
Expenses:
F=C*30% Payment to closed circuit theatre managers £          4,290,000
G Payment to Bands £          7,000,000
H=F+G Total Expenses £           11,290,000
I=E-H Most BBT Productions can spend Advertising £             3,310,000
2) A Required Operating Income £              500,000
B Total fixed Costs £          9,000,000
C Less:Receipt from London Arena £              300,000
D=A+B-C Total Contribution required £          9,200,000
E Sales Price per ticket £                   13.00
F=E*30% Payment to closed circuit theatre managers per ticket £                     3.90
G=E-F Contribution per ticket £                     9.10
N=D/G Number of tickets required to be sold               1,010,989
Operating Income Calculation
A Number of tickets sold               1,010,989
B Collection per ticket £                   13.00
C=A*B Total collection from ticke sale £        13,142,857
D Receipt from London Arena £              300,000
E=C+D Total Receipts £           13,442,857
Expenses:
F=C*30% Payment to closed circuit theatre managers £          3,942,857
G Fixed expenses £          9,000,000
H=F+G Total Expenses £           12,942,857
I=E-H Operating Income £                 500,000